OpenStudy (anonymous):

How do I solve H squared minus 4H plus 1 =0?

OpenStudy (anonymous):

\[h^2-4h+1 = 0\]use the quadratic formula to solve this, if you have an equation in the form:\[ax^2+bx+c=0\]you have find the values for x (also called roots) with the followng formula:\[x = {-b \pm \sqrt{b^2-4ac} \over 2a} \] did this help?